OpenStudy (anonymous):

What is the degree measure of an arc 4 ft. long in a circle of radius 10 ft

OpenStudy (anonymous):

Wanna Put Medal For the Question... how to put that?

OpenStudy (anonymous):

Find the Circumference of the circlle C=2*pi*r C=2*pi*10=20*pi create a proportion 20pi is to 360 degrees as 4pi is to x degrees 20*pi/360=4*pi/x solve for x x=360*4*pi/(20*pi)=72 degrees

OpenStudy (anonymous):

\[\frac{ arclength }{ circumference }=\frac{ degreemeasure }{ 360 }\] in simpler term @malachip